# Attention Sinks in Diffusion Language Models

### v1

@@ 1 Introduction
Large Language Models (LLMs) have driven a paradigm shift across numerous scientific and industrial domains, demonstrating remarkable capabilities in language understanding, generation, and reasoning (Achiam et al., 2023; Anthropic, 2025; Yang et al., 2025a; MetaAI, 2024). This rapid progress is rooted in the transformer architecture and the attention mechanism (Vaswani et al., 2017). While attention is a critical aspect of the transformer’s effectiveness, it also gives rise to complex and often non-intuitive emergent phenomena.
One of the most striking traits of these behaviours is the "attention sink" (Xiao et al., 2023; Miller, 2023). This consists in the fact that, in most autoregressive models (ARMs), a small subset of tokens consistently receives a disproportionate amount of attention from other tokens in the sequence. The pattern is not limited to language, and similar patterns have been observed in Vision Transformers (Darcet et al., 2024) and encoder-only transformers (Ruscio et al., 2025), suggesting it may be a fundamental property of attention-based deep networks.
Figure 1: Incoming attention scores for each token in LLaDA-8B (Nie et al., 2025) across denoising steps. Unlike autoregressive models, DLMs exhibit attention sinks that shift across the sequence as tokens are progressively unmasked.
Recently, masked Diffusion Language Models (DLMs) have emerged as an alternative to the dominant autoregressive paradigm (Nie et al., 2025; Ye et al., 2025c; Team, 2025; Labs et al., 2025; Yang et al., 2025b; Wang et al., 2025; Song et al., 2025; Sahoo et al., 2024; Zhu et al., 2025; Liu et al., 2025). Unlike Autoregressive Models (ARMs), which generate text strictly from left to right, DLMs iteratively refine a fully masked sequence through successive denoising steps (Nie et al., 2025; Ye et al., 2025c; Yang et al., 2025b). Generation is based on the unmasking of an initial fully masked sequence of tokens, that the model progressively "denoises" over multiple steps to produce a coherent fully unmasked output. Crucially, DLMs employ a bidirectional attention mechanism. While this bidirectional information flow is key to their parallel, non-causal generation process, the precise impact of this architecture on the inner workings of DLMs remains largely unexplored.
In this work, we present an empirical study of attention patterns in DLMs, focusing specifically on the attention sink phenomenon. We analyse three state-of-the-art open-source masked DLMs: Dream-7B (Ye et al., 2025c), a model initialized from a pre-trained ARM; LLaDA-8B (Nie et al., 2025), a large-scale model trained from scratch; and MMaDA-8B (Yang et al., 2025b), a multimodal DLM trained from LLaDA-8B. Our analysis reveals that DLMs do exhibit attention sinks, but these sinks possess unique dynamic properties rarely seen in their autoregressive counterparts.
@@ 2.1 Diffusion Language Models
Language modelling has traditionally been dominated by autoregressive models that generate text sequentially, one token at a time. While this paradigm has proven highly successful, DLMs have emerged as an alternative, offering token generation through an iterative denoising processes with potential efficiency advantages (Li et al., 2025b; Wu et al., 2025b; Kim et al., 2025; Liu et al., 2025; Li et al., 2025a; Wu et al., 2025a). Some applications of diffusion to language modelling operate in continuous space, first embedding discrete tokens into continuous vectors, applying diffusion-based denoising, and then mapping back to discrete tokens (Li et al., 2022; Strudel et al., 2022; Gong et al., 2022; Dieleman et al., 2022).
While theoretically elegant, this approach introduces additional complexity in handling the discrete nature of language. A more direct approach emerged with discrete diffusion models, which operate directly on token vocabularies (Austin et al., 2021; Gong et al., 2023; Hoogeboom et al., 2021; Campbell et al., 2022). Starting from fully masked sequences of [MASK] tokens, these models iteratively predict and refine tokens through a process reminiscent of BERT-style masked language modelling (He et al., 2023; Gong et al., 2025). Several works (Austin et al., 2021; He et al., 2023; Gong et al., 2025) have adopted this paradigm but faced significant scaling challenges, remaining limited in size while autoregressive models scaled to billions of parameters.
Recently, discrete DLMs have gained traction thanks to open-source models like Dream-7B (Ye et al., 2025c), MMaDA-8B (Yang et al., 2025b) and LLaDA-8B (Nie et al., 2025; Zhu et al., 2025; Liu et al., 2025), which have successfully scaled to 7 billion parameters and beyond, narrowing the performance gap with ARMs.
@@ 2.2 Attention Sink in Transformers
Attention Sink refers to the common phenomenon observed in transformers where a small subset of tokens consistently receives a disproportionate amount of attention from other tokens in the sequence. This behaviour was initially discovered in Xiao et al. (2023), and leveraged for efficiency. After this, other works have then explored the sink phenomenon, characterizing properties of sink tokens like high $L_{2}$ norm in the hidden state activations (Sun et al., 2024; Cancedda, 2024) or low $L_{2}$ norm in the key projection (Devoto et al., 2024; Gu et al., 2024). Similar properties have been also observed in the vision domain (Darcet et al., 2024). Several works have attempted to explain the emergence of attention sinks in transformers. Gu et al. (2025) offers an empirical study of how attention sinks manifest in transformer models, specifically focusing on ARMs. Barbero et al. (2024, 2025) and Pappone (2025) investigate the phenomenon analytically and show how attention sinks act as a bias for ARMs and can mitigate information over-squashing. Finally,  Ruscio et al. (2025) analyses attention sinks from a geometric perspective, and shows that they emerge to establish stable coordinate systems in the model’s high-dimensional latent space. While these works analyse sinks in both decoder and encoder transformers, we are the first to observe and investigate this phenomenon in the context of DLMs.
@@ 3 Background on Masked Discrete Diffusion
Traditional ARMs model the probability of a text sequence $\mathbf{x}=(x_{1},x_{2},\dots,x_{L})$ of length $L$ by decomposing the joint probability into a product of conditional probabilities, generated in a strict, left-to-right order (Jelinek, 1980; Bengio et al., 2000). This decomposition is given by:
@@
In this work we consider three masked discrete DLMs: LLaDA-8B (Nie et al., 2025), MMaDA-8B (Yang et al., 2025b) and Dream-7B (Ye et al., 2025c). LLaDA-8B and MMaDA-8B are trained from scratch, with a masked language modelling loss where a token $x_{i}$ is masked during the forward process, and the model learns to predict the token itself ($x_{i}\rightarrow$ [MASK] $\rightarrow x_{i}$). At inference time, LLaDA-8B and MMaDA-8B use semi-autoregressive block diffusion, where the input sequence is divided into blocks, and the model gradually unmasks all tokens inside the corresponding block in a left-to-right manner (Arriola et al., 2025), (see Figure˜2). Dream-7B, on the other hand, is initialized from an autoregressive model to leverage the pretrained weights and its training objective employs a "shift operation" (Ye et al., 2025c; Gong et al., 2025). More specifically, when a token $x_{i}$ is masked, Dream-7B is trained to predict $x_{i+1}$, similarly to an autoregressive model ($x_{i}\rightarrow$ [MASK] $\rightarrow x_{i+1}$). In Figure˜2 we provide a comparison and visual explanation of how the different types of inference are implemented.
@@ 4 Analysis of Attention Sinks in Masked Diffusion Language Models
Previous work has shown that attention sinks emerge in most transformer-based architectures, regardless of the data domain and training strategy (Gu et al., 2025; Ruscio et al., 2025; Xiao et al., 2023; Darcet et al., 2024). Attention sinks are characterized by the disproportionate attention score they receive from all the tokens in the sequence, and can be easily identified as vertical bright lines in attention maps (like the one we show in Figure˜1). To validate the presence of attention sinks in DLMs, we first analyse the distribution of attention scores in LLaDA-8B and show it in Figure˜3. We see that only a few tokens, the sinks, capture a very high attention score consistently. Similar patterns emerge for Dream-7B and MMaDA-8B (see Appendix˜A). We now define a metric to characterize and locate attention sinks in DLMs.
Figure 3: Distribution of attention scores in LLaDA-8B (Nie et al., 2025) across denoising steps. Only a few tokens, the attention sinks, receive a very high attention score, while the majority of tokens in the sequence have scores close to zero.
@@ 4.1 Definition of Attention Sink
$j\text{ is a sink token if }\bar{A}^{(l,h)}_{j}>\frac{1}{S-1}\sum_{k\neq j}\bar{A}^{(l,h)}_{l}+\epsilon$ (3)
@@ 4.2 Sink Patterns
showcases a sink behaviour that follows primarily a positional rather than a semantic pattern. Unlike LLaDA-8B, Dream-7B’s sinks often originate at the rightmost masked token and shift leftward as tokens are progressively unmasked, regardless of the token content, as we show in Figure˜8(b). This right-to-left migration is most prominent in early layers and creates a dynamic attention flow that follows the unmasking frontier. This positional nature of Dream-7B’s sinks likely stems from its initialization from a pre-trained autoregressive model. The inherited representations may be less refined for bidirectional attention, causing the model to rely on positional cues rather than semantic content for sink formation. Dream-7B’s positional bias represents a difference from LLaDA-8B’s semantic approach and suggests that initialization strategy and positional embeddings significantly influences attention organization in diffusion models (Ruscio et al., 2025).
@@
In Figure˜7 we show how sinks behave in different models. We select a specific head from each model and compare the position of the largest sink detected by our metric. We observe that while MMaDA-8B exhibits a mostly static sinking behaviour, sinks tend to shift position in Dream-7B and LLaDA-8B. More specifically, we observe that in LLaDA-8B the sink tends to shift right as more blocks are denoised, while it moves from right to left in Dream-7B.
Dataset Sinks Dream-7B [58] LLaDA-8B [35] MMaDA-8B [54] Llama-3.1-8B [33]
@@ 4.3 Robustness of DLMs to Masking Sinks
Previous studies have demonstrated that attention sinks play a crucial role in transformer-based models, with their removal typically causing catastrophic performance degradation (Xiao et al., 2023; Gu et al., 2025; Barbero et al., 2024). However, given that attention sinks in DLMs exhibit markedly different and more dynamic patterns compared to ARMs, we investigate whether DLMs demonstrate similar sensitivity to sink masking during generation.
We evaluate the three DLM variants — LLaDA-8B, Dream-7B, and MMaDA-8B — on both coding and mathematical reasoning tasks using the GSM8K (Cobbe et al., 2021) and HumanEval Chen et al. (2021) datasets. GSM8K contains grade-school level math word problems, while HumanEval comprises programming problems designed to evaluate code generation and reasoning capabilities. For each model, we conduct two sets of evaluations: (1) using the original, unmodified model, and (2) masking attention scores directed toward the top-K attention sinks identified by our metric (Equation 3). We vary the threshold parameter $\epsilon$, where smaller values result in masking a larger proportion of sinks. Specifically, we select $\epsilon_{0}$, $\epsilon_{1}$ and $\epsilon_{2}$ to mask the top $1,5\text{ and }10$ sinks respectively.
Surprisingly, the tested DLMs exhibit only modest performance degradation when sinks are masked (Table 1). For all the tested DLMs, masking one sink leads to a degradation in performance smaller than 1%. Substantial degradation occurs only when $\epsilon$ is decreased further to mask $10$ sinks, and mostly in MMaDA-8B. In contrast, applying the same masking procedure to LLama-3.1-8B results in severe performance drops even when masking a single sink token, confirming prior findings that ARMs are highly sensitive to attention sink removal (Xiao et al., 2023; Gu et al., 2025).
@@
We evaluate our models in PyTorch (Paszke et al., 2019) using the checkpoints released on Hugging Face transformers (Wolf et al., 2020) and the official lm evaluation harness scripts (Gao et al., 2024). We use the same hyper-parameters specified in the respective original papers. For LLaDA-8B, we use a block size of 32 and a generation length of 256 tokens for GSM8K and 512 for HumanEval. For Dream-7B, which does not use semi-autoregressive block generation, we adjust only the generation length and diffusion step parameters according to the original settings. We successfully reproduce the reported results for LLaDA-8B, Dream-7B, and LLama-3.1-8B using these configurations. However, we were unable to reproduce the original results for MMaDA-8B despite following the published implementation details, and we therefore report our own evaluation results for this model. Throughout our analysis, we employ $\epsilon=3$ for sink detection, a threshold that empirically balances robust sink identification with the exclusion of non-sink tokens.
@@ 5.1 Dynamic Sinks and Positional Encoding
Recent work on encoder-only models notes that attention sinks can shift usually around special markers like [CLS] or [EOS] and connects this behaviour to the use of absolute positional embeddings (Ruscio et al., 2025). However, we find that DLMs, despite using Rotary Positional Embeddings (RoPE,  Su et al. 2023), show extremely varied and dynamic sink patterns, including sinks that move and others that split attention between masked and unmasked tokens. These appear all over the text sequence, often on important structural tokens (like punctuation). The emergence of sink tokens on semantic markers suggests that the sinking behaviour is driven not only by the positional encoding or token index in the sequence (Ruscio et al., 2025; Barbero et al., 2025), but also by training dynamics and frequency of the token in the training corpus Sun et al. (2024); Land and Bartolo (2024).

### v2

@@ 1 Introduction
Correspondence to: rulli.2154435@studenti.uniroma1.it, alessio.devoto@uniroma1.it
Large Language Models (LLMs) have driven a paradigm shift across numerous scientific and industrial domains, demonstrating remarkable capabilities in language understanding, generation, and reasoning (gpt4; claude; qwen3; llama4). This rapid progress is rooted in the transformer architecture and the attention mechanism (transformer). While attention is a critical aspect of the transformer’s effectiveness, it also gives rise to complex and often non-intuitive emergent phenomena.
One of the most striking traits of these behaviours is the "attention sink" (streamingllm; offbyone). This consists in the fact that, in most autoregressive models (ARMs), a small subset of tokens consistently receives a disproportionate amount of attention from other tokens in the sequence. The pattern is not limited to language, and similar patterns have been observed in Vision Transformers (registers) and encoder-only transformers (rusciogeometric), suggesting it may be a fundamental property of attention-based deep networks.
Figure 1: Incoming attention scores for each token in LLaDA-8B (llada) across denoising steps. Unlike autoregressive models, DLMs exhibit attention sinks that shift across the sequence as tokens are progressively unmasked.
Recently, masked Diffusion Language Models (DLMs) have emerged as an alternative to the dominant autoregressive paradigm (llada; dream; geminidiff; inceptiondiffusion; mmada; wangrevolutionizing; songseed; sahoosimple; lladamoe; longllada). Unlike Autoregressive Models (ARMs), which generate text strictly from left to right, DLMs iteratively refine a fully masked sequence through successive denoising steps (llada; dream; mmada). Generation is based on the unmasking of an initial fully masked sequence of tokens, that the model progressively "denoises" over multiple steps to produce a coherent fully unmasked output. Crucially, DLMs employ a bidirectional attention mechanism. While this bidirectional information flow is key to their parallel, non-causal generation process, the precise impact of this architecture on the inner workings of DLMs remains largely unexplored.
In this work, we present an empirical study of attention patterns in DLMs, focusing specifically on the attention sink phenomenon. We analyse three state-of-the-art open-source masked DLMs: Dream-7B (dream), a model initialized from a pre-trained ARM; LLaDA-8B (llada), a large-scale model trained from scratch; and MMaDA-8B (mmada), a multimodal DLM trained from LLaDA-8B. Our analysis reveals that DLMs do exhibit attention sinks, but these sinks possess unique dynamic properties rarely seen in their autoregressive counterparts.
@@ 2.1 Diffusion Language Models
Language modelling has traditionally been dominated by autoregressive models that generate text sequentially, one token at a time. While this paradigm has proven highly successful, DLMs have emerged as an alternative, offering token generation through an iterative denoising processes with potential efficiency advantages (dlmssurvey; wufast; kimany; longllada; libeyond; wufast-2). Some applications of diffusion to language modelling operate in continuous space, first embedding discrete tokens into continuous vectors, applying diffusion-based denoising, and then mapping back to discrete tokens (li2022diffusion; strudel2022self; gongdiffuseq; dieleman2022continuous).
While theoretically elegant, this approach introduces additional complexity in handling the discrete nature of language. A more direct approach emerged with discrete diffusion models, which operate directly on token vocabularies (austin2021structured; gongscaling; hoogeboom2021argmax; campbell2022continuous). Starting from fully masked sequences of [MASK] tokens, these models iteratively predict and refine tokens through a process reminiscent of BERT-style masked language modelling (he2023diffusionbert; gong2025scalingdiffusionlanguagemodels). Several works (austin2021structured; he2023diffusionbert; gong2025scalingdiffusionlanguagemodels) have adopted this paradigm but faced significant scaling challenges, remaining limited in size while autoregressive models scaled to billions of parameters.
Recently, discrete DLMs have gained traction thanks to open-source models like Dream-7B (dream), MMaDA-8B (mmada) and LLaDA-8B (llada; lladamoe; longllada), which have successfully scaled to 7 billion parameters and beyond, narrowing the performance gap with ARMs.
@@ 2.2 Attention Sink in Transformers
Attention Sink refers to the common phenomenon observed in transformers where a small subset of tokens consistently receives a disproportionate amount of attention from other tokens in the sequence. This behaviour was initially discovered in streamingllm, and leveraged for efficiency. After this, other works have then explored the sink phenomenon, characterizing properties of sink tokens like high $L_{2}$ norm in the hidden state activations (massive; cancedda) or low $L_{2}$ norm in the key projection (knorm; guattention). Similar properties have been also observed in the vision domain (registers). Several works have attempted to explain the emergence of attention sinks in transformers. when offers an empirical study of how attention sinks manifest in transformer models, specifically focusing on ARMs. glasses; barbero2025llmsattendtoken and pappone2025attentionsinks investigate the phenomenon analytically and show how attention sinks act as a bias for ARMs and can mitigate information over-squashing. Finally,  rusciogeometric analyses attention sinks from a geometric perspective, and shows that they emerge to establish stable coordinate systems in the model’s high-dimensional latent space. While these works analyse sinks in both decoder and encoder transformers, we are the first to observe and investigate this phenomenon in the context of DLMs.
@@ 3 Background on Masked Discrete Diffusion
Traditional ARMs model the probability of a text sequence $\mathbf{x}=(x_{1},x_{2},\dots,x_{L})$ of length $L$ by decomposing the joint probability into a product of conditional probabilities, generated in a strict, left-to-right order (jelinek1980interpolated; bengio2000neural). This decomposition is given by:
@@
In this work we consider three masked discrete DLMs: LLaDA-8B (llada), MMaDA-8B (mmada) and Dream-7B (dream). LLaDA-8B and MMaDA-8B are trained from scratch, with a masked language modelling loss where a token $x_{i}$ is masked during the forward process, and the model learns to predict the token itself ($x_{i}\rightarrow$ [MASK] $\rightarrow x_{i}$). At inference time, LLaDA-8B and MMaDA-8B use semi-autoregressive block diffusion, where the input sequence is divided into blocks, and the model gradually unmasks all tokens inside the corresponding block in a left-to-right manner (arriolablock), (see Figure˜2). Dream-7B, on the other hand, is initialized from an autoregressive model to leverage the pretrained weights and its training objective employs a "shift operation" (dream; gong2025scalingdiffusionlanguagemodels). More specifically, when a token $x_{i}$ is masked, Dream-7B is trained to predict $x_{i+1}$, similarly to an autoregressive model ($x_{i}\rightarrow$ [MASK] $\rightarrow x_{i+1}$). In Figure˜2 we provide a comparison and visual explanation of how the different types of inference are implemented.
@@ 4 Analysis of Attention Sinks in Masked Diffusion Language Models
Previous work has shown that attention sinks emerge in most transformer-based architectures, regardless of the data domain and training strategy (when; rusciogeometric; streamingllm; registers). Attention sinks are characterized by the disproportionate attention score they receive from all the tokens in the sequence, and can be easily identified as vertical bright lines in attention maps (like the one we show in Figure˜1). To validate the presence of attention sinks in DLMs, we first analyse the distribution of attention scores in LLaDA-8B and show it in Figure˜3. We see that only a few tokens, the sinks, capture a very high attention score consistently. Similar patterns emerge for Dream-7B and MMaDA-8B (see Appendix˜A). We now define a metric to characterize and locate attention sinks in DLMs.
Figure 3: Distribution of attention scores in LLaDA-8B (llada) across denoising steps. Only a few tokens, the attention sinks, receive a very high attention score, while the majority of tokens in the sequence have scores close to zero.
@@ 4.1 Definition of Attention Sink
$j\text{ is a sink token if }\bar{A}^{(t,l,h)}_{j}>\frac{1}{S-1}\sum_{k\neq j}\bar{A}^{(t,l,h)}_{k}+\epsilon$ (3)
@@ 4.2 Sink Patterns
showcases a sink behaviour that follows primarily a positional rather than a semantic pattern. Unlike LLaDA-8B, Dream-7B’s sinks often originate at the rightmost masked token and shift leftward as tokens are progressively unmasked, regardless of the token content, as we show in Figure˜8(b). This right-to-left migration is most prominent in early layers and creates a dynamic attention flow that follows the unmasking frontier. This positional nature of Dream-7B’s sinks likely stems from its initialization from a pre-trained autoregressive model. The inherited representations may be less refined for bidirectional attention, causing the model to rely on positional cues rather than semantic content for sink formation. Dream-7B’s positional bias represents a difference from LLaDA-8B’s semantic approach and suggests that initialization strategy and positional embeddings significantly influences attention organization in diffusion models (rusciogeometric).
@@
In Figure˜7 we show an example of the phenomena described previously for the different models. We select a specific head from each model and compare the position of the largest sink detected by our metric. We observe that while MMaDA-8B exhibits a mostly static sinking behaviour, sinks tend to shift position in Dream-7B and LLaDA-8B. More specifically, we observe that in LLaDA-8B the sink tends to shift right as more blocks are denoised, while it moves from right to left in Dream-7B.
Dataset Sinks Dream-7B [dream] LLaDA-8B [llada] MMaDA-8B [mmada] Llama-3.1-8B [llama3]
@@ 4.3 Robustness of DLMs to Masking Sinks
Previous studies have demonstrated that attention sinks play a crucial role in transformer-based models, with their removal typically causing catastrophic performance degradation (streamingllm; when; glasses). However, given that attention sinks in DLMs exhibit markedly different and more dynamic patterns compared to ARMs, we investigate whether DLMs demonstrate similar sensitivity to sink masking during generation.
We evaluate the three DLM variants — LLaDA-8B, Dream-7B, and MMaDA-8B — on both coding and mathematical reasoning tasks using the GSM8K (gsm8k) and HumanEval humaneval datasets. GSM8K contains grade-school level math word problems, while HumanEval comprises programming problems designed to evaluate code generation and reasoning capabilities. We evaluate each model in two configurations: (1) the original, unmodified model, and (2) with the top-K sink tokens masked. To determine which tokens to mask, we analyse the attention map from step t1t-1 across each head and layer. We identify the top-K sink tokens based on our metric (Equation 3) and mask them by setting their attention to -$\infty$ during the computation of step $t$. We vary the threshold parameter $\epsilon$, where smaller values result in masking a larger proportion of sinks. Specifically, we select $\epsilon_{0}$, $\epsilon_{1}$ and $\epsilon_{2}$ to mask the top $1,5\text{ and }10$ sinks respectively.
Surprisingly, the tested DLMs exhibit only modest performance degradation when sinks are masked (Table 1). For all the tested DLMs, masking one sink leads to a degradation in performance smaller than 1%. Substantial degradation occurs only when $\epsilon$ is decreased further to mask $10$ sinks, and mostly in MMaDA-8B. To make a fair comparison with ARMs, we conduct a similar experiment with LLama-3.1-8B, where instead of analyzing the entire attention map to determine the token to mask at the next iteration, we focus on the attention distribution of the last token. We adopt this approach because ARMs, due to their causal nature, cannot ’shift’ their attention as flexibly as DLMs. In contrast to DLMs, applying this masking procedure to LLama-3.1-8B results in severe performance drops even when masking a single sink token, confirming prior findings that ARMs are highly sensitive to attention sink removal (streamingllm; when).
@@
We evaluate our models in PyTorch (torch) using the checkpoints released on Hugging Face transformers (transformers) and the official lm evaluation harness scripts (eval-harness). We use the same hyper-parameters specified in the respective original papers. For LLaDA-8B, we use a block size of 32 and a generation length of 256 tokens for GSM8K and 512 for HumanEval. For Dream-7B, which does not use semi-autoregressive block generation, we adjust only the generation length and diffusion step parameters according to the original settings. We successfully reproduce the reported results for LLaDA-8B, Dream-7B, and LLama-3.1-8B using these configurations. However, we were unable to reproduce the original results for MMaDA-8B despite following the published implementation details, and we therefore report our own evaluation results for this model. Throughout our analysis, we employ $\epsilon=3$ for sink detection, a threshold that empirically balances robust sink identification with the exclusion of non-sink tokens.
@@ 5.1 Dynamic Sinks and Positional Encoding
Recent work on encoder-only models notes that attention sinks can shift usually around special markers like [CLS] or [EOS] and connects this behaviour to the use of absolute positional embeddings (rusciogeometric). However, we find that DLMs, despite using Rotary Positional Embeddings (RoPE,  rope), show extremely varied and dynamic sink patterns, including sinks that move and others that split attention between masked and unmasked tokens. These appear all over the text sequence, often on important structural tokens (like punctuation). We show relative frequencies of sink tokens in Table 2. The emergence of sink tokens on semantic and formatting markers suggests that the sinking behaviour is driven not only by the positional encoding or token index in the sequence (rusciogeometric; barbero2025llmsattendtoken), but also by training dynamics and frequency of the token in the training corpus massive; magikarp.